OpenStudy (anonymous):

Dad is 4 times as old as his son Jim. In 10 years, Dad's age will be 20 years more than twice Jim's age. How old is Jim? 10 years old 15 years old 30 years old 10??

OpenStudy (anonymous):

D = 4J D+10 = 2(10 + J) + 20 D = 30 + 2J 30 + 2J = 4J J = 15

OpenStudy (anonymous):

Dad`s age now = y Jim`s age now = x Dad is 4 times as old as his son Jim so: 4x=y In ten years, Dads age will be twenty years more than twice Jims age. then y+10=2(x+10)+20 then 4x+10=2x+40 2x=30 x=15 and y=60

OpenStudy (anonymous):

do you understand?

OpenStudy (anonymous):

yesh thanks for explaing it

OpenStudy (anonymous):

you Are welcome. Any more questions

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i should have more in the future thanks

OpenStudy (anonymous):

Ok. Happy to Help :D